Determine the nature of function y=f(I) which equals demand for money with supply of money Ms=3950 Md=L1+ L2 L1=0.7y L2=1850-77i. Y=4210.35-100i
1
Expert's answer
2021-10-27T09:16:34-0400

"M_d=L_1+L_2"

"M_d=0.7y+1850-77i"

"M_s=3950"

"Y=4210.35-100i"

"Y=M_d+M_s"

"Y=0.7y+1850-77i+3950"

"0.7y+5800-77i=4210.35-100i"

"0.7y=-1589.65-23i"

"y=\frac{-1589.65-23i}{0.7}"